OpenStudy (anonymous):

Vector Calculus Problem: I have to find where these two planes intersect. I know this should be really easy for me but for some reason i'm not getting it 3x-2y+z=10 and -4x+5y+2z=14

OpenStudy (amistre64):

lets see where they hit the xy plane when z=0 3x-2y=10 -4x+5y=14 y = 82/7 x = 78/7 lets see where they cross when z=....10

OpenStudy (amistre64):

3x-2y+10=10 3x -2y = 0 -4x+5y=-6 y= -18/7 x = -12/7

OpenStudy (amistre64):

so we have 2 points now of intersection I believe; lets line them up :)

OpenStudy (amistre64):

( 78/7, 82/7, 0) -(-12/7,-18/7,10) ----------------- < 90/7,100/7,-10>

OpenStudy (amistre64):

t< 9,10,-1/7> x=78x+9t y=82y+10t z=-(1/7)t maybe?

OpenStudy (amistre64):

the wolf might suggest its t<9,10,-7>

OpenStudy (amistre64):

<9/15, 10/15,-7/15> <9,10,-7> = vector parallel to line